# Predictive Models in Sequential Recommendations: Bridging Performance Laws with Data Quality Insights

### v3

@@ 3.3 Theorem of Performance Law
Assuming that the user sequence can be modeled as a first-order aperiodic stationary Markov chain. If the user sequence is $S=\{S_{u},u\in U\}$, then the minimum encoding length is given by:
$L(C)\geq(\Sigma_{i=1}^{n}|s_{i}|)/ApEn(S)$ (12)
@@
$n\sum_{\forall s_{i}}p(s_{i})\log_{2}\frac{1}{p(s_{i})}\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})},$ (15)
@@
$\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{nS_{max}},$ (17)
@@
The right-hand side is non-zero only when it appears in the dataset, and each term has a minimum value of $\frac{1}{n\cdot S_{max}}$. Subsequently, we have:
$\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})\geq\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{nS_{max}}.$ (20)

### v4

@@ 3.3 Theorem of Performance Law
With the aforementioned lemma established, we proceed to conduct a detailed analysis of data volume within the framework of the original Scaling Law. Importantly, the minimum encoding length L(C) provides a nuanced perspective by integrating data compression rates (sayood2017introduction; Yin et al., 2024b), thereby offering insights into both the scale and quality of the data. This refined approach permits a more thorough assessment of how data affects model performance. Guided by these insights, we propose the following theorem:
@@
Assuming that the user sequence can be modeled as a first-order aperiodic stationary Markov chain. If the user sequence is $S=\{S_{u},u\in U\}$, then the summary of minimum encoding lengths is given by:
$|U|L(C)\geq(\Sigma_{i=1}^{n}|s_{i}|)/ApEn(S)$ (12)
@@
$|U|\sum_{\forall s_{i}}p(s_{i})\log_{2}\frac{1}{p(s_{i})}\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})},$ (15)
@@
$\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{|U|S_{max}},$ (17)
@@
The right-hand side is non-zero only when it appears in the dataset, and each term has a minimum value of $\frac{1}{|U|\cdot S_{max}}$. Subsequently, we have:
$\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}log(\frac{1}{\mathrm{p_{xy}}})\geq\sum_{\mathrm{x\in I}}\sum_{\mathrm{y\in I}}\pi(\mathrm{x})\mathrm{p_{xy}}\geq\frac{\Sigma_{i=1}^{n}|s_{i}|}{|U|S_{max}}.$ (20)